tickets for the senior play cost $4.00 for adults and $2.00 for students. this year there were 600 tickets sold, and the class made $1900. how many of each type ticket were sold?
Answer by angelmidnight(6)

4x+2y=1900
-2x-2y=-1200
2x=700
x=350
4*350+2y=1900
1400+2y=1900
2y=500
y=250
They sold 350 Adult tickets and 250 Student tickets.
4*350+2*250=1900